# Modified Abelian and SU(2) Wilson theories on a lattice from a non-compact regularization

### v1

@@ Abstract
Multiflavour gauge theories of matter systems on a three-dimensional lattice have recently been widely investigated especially in connection with a possible symmetry enlargement at a continuous phase transition. Abelian models were studied both with compact gauge fields and in a mixed formulation in which the coupling with matter fields is in compact form while the gauge fields Lagrangian is written in terms of non-compact gauge fields, getting quite different results. Such a mixed formulation is not permissible for non-Abelian theories, for which however there exists an entirely non-compact formulation (in which exact gauge invariance is enforced by help of auxiliary fields), that for SU(2) was shown to yield in the scaling window a larger physical volume than Wilson’s one. The corresponding U(1) non-compact regularization is derived in the present work. In both Abelian and SU(2) cases there is only one auxiliary field which for large mass has a linear coupling with the other fields and it can be integrated out. The coupling with the auxiliary field might make the inversion of the fermion matrix easier.
@@ I Introduction
The initial motivation of the present study was the recent interest in multiflavor gauge theories in three space-time dimensions [Peli1, ; Bona1, ; higher-charge, ; Bona2, ], but we think that the results have a wider perspective. Multiflavor gauge theories appear in the investigation of condensed matter [Klei, ] and critical phenomena, especially in connection with a possible symmetry enlargement at a second order phase transition [Sach, ]. The relative non-perturbative investigations [Peli1, ; Bona1, ; higher-charge, ; Bona2, ] have been mostly performed by lattice Monte Carlo simulations and in several Abelian models the phase diagrams turned out to depend on the way, compact or non-compact, the gauge fields are defined on a lattice.
@@
Later on a new regularization of GL($N_{c}$) theories on a lattice was defined [Palu1, ], in which the gauge fields are maintained in the algebra of the group, namely they are non-compact, and in order to enforce gauge invariance auxiliary fields are introduced which decouple in the continuum limit. Since also the auxiliary fields are non-compact, such a regularization can be called non-compact. For $N_{c}=2$ the irreducible representation has an SU(2) invariance [Becc1, ], while the general definition for SU($N_{c}>2$) symmetry has been given in a subsequent work [Palu2, ]. The construction of such non-compact regularization was done with the idea of being closer to the continuum at finite lattice spacing, an expectation supported by a Monte Carlo simulation in four dimensions which showed that in the scaling window it generates for SU(2) a physical space larger than the Wilson’s one [DiCa, ].
@@
In the study of Abelian theories another regularization [Bona1, ] is used, in which the gauge fields Lagrangian is written as a discretization of the continuous one, namely it is non-compact, while the coupling to matter fields is written in terms of Wilson variables. For this reason, and to distinguish it from the entirely non-compact regularization [Palu1, ] that we will use, we will call it half-compact, even though in the literature it is generally referred to as non-compact, hoping that this will not generate any confusion.
@@
Compact and half-compact regularizations in three dimensions give quite different phase diagrams. In the first case the flavor symmetry is broken while the gauge fields are in the confined phase; in the second case there are three phases: the Higgs and molecular phases, in which the flavor symmetry is broken and the electromagnetic correlations are gapped and ungapped, respectively, and the Coulomb phase in which the flavor symmetry is unbroken and the electromagnetic correlations are long range. In view of such striking difference it seems worth while to compare them with the entirely non-compact one [Palu1, ] which might also introduce further physical behaviors.
@@
The half-compact regularization cannot be used for non-Abelian theories, because it breaks the gauge invariance of gauge fields self-interactions and therefore a comparison between compact and half-compact was restricted to the Abelian case. Also for non-Abelian theories, however, a comparison can be made with the non-compact regularization, for which there is the additional expectation of a faster convergence to the continuum.
@@
As a first approach instead of studying entirely non-compact models, one might investigate the corrections they give to the Wilson Lagrangian at small lattice spacing, in analogy to the modifications obtained by adding some extra terms chosen to cancel the leading lattice artifacts or using the Renormalization Group transformations [Syma, ; Nied, ]. It is the purpose of the present work to derive such corrections.
@@
SU(2) and U(1) theories, which are the ones investigated in the quoted multiflavour models [Peli1, ; Bona1, ; higher-charge, ; Bona2, ] have in their non-compact form one feature in common: they have only one auxiliary field which in polar variables is a gauge invariant function of the sites. For large mass this field has linear couplings with all the other (gauge and matter) fields. The resulting Lagrangian is analogous to the gauged Nambu–Jona-Lasinio model [Kogu2, ] after linearization: the auxiliary field can then be integrated out leaving a Lagrangian expressed in terms of Wilson variables and matter fields only. Of course, in the presence of fermion fields the integration over the auxiliary field has to be performed after that on the Grassmann variables, but in analogy with the gauged Nambu–Jona-Lasinio model, the presence of the auxiliary field might make easier the inversion of the fermion matrix (but see later on this point).
@@
The half-compact and non-compact regularizations differ from each other also in another point: only the first one needs a gauge fixing to get rid of the pure gauge degrees of freedom which have an infinite volume in the partition function. Some gauge fixings, however, are not compatible with some boundary conditions (see for instance Ref. [Bona3, ]),while the choice of boundary conditions may have physical consequences, a point illustrated by a few examples in Appendix A. In the non-compact regularization [Palu1, ] instead no gauge fixing is necessary because the mentioned gauge invariant potential damps all the gauge fields, physical and auxiliary, so that all boundary conditions can be handled on the same footing.
@@
Abelian theories have been compared in the compact and half-compact forms. It was soon established that the first one does not yield a second order phase transition in the bare coupling constant and therefore it cannot define a continuum theory [Kogu1, ]. Several investigations have been performed afterwards using the half-compact formulation. After the result that at small coupling the theory is trivial also in this case, it was searched whether a continuum theory could be defined at finite coupling [Azco1, ], until it was established that models of Abelian gauge fields coupled to scalars [Baig, ] or fermions with four flavours in the chiral limit [Kogu2, ] are indeed trivial. Later on, however, evidence was reported that in a large-$N$ expansion an ultraviolet fixed point exists in a theory of gauge fields coupled to both fermions and scalars [Liti, ].
@@
Non-Abelian theories instead have been studied in compact and non-compact form, the latter one however restricted to SU(2) and only in the absence of matter fields. Firstly it was established that the non-compact formulation gives an evidence for confinement [Poli1, ] comparable to that of Wilson. This was a prerequisite, because non-compact gauge fields with a simple discretization of the continuum theory [Patr, ] or a non-renormalizable gauge fixing [Poli2, ] gave a vanishing string tension. Next a perturbative study of the scaling properties determined the renormalization scale parameter and showed that the theory has the same fixed point as Wilson’s one [Becc2, ]. This result was confirmed by a calculation [Bora, ] in the Hamiltonian formalism [Dick, ] using Lüscher’s effective Hamiltonian [Lusc, ]. Lastly, as already said, the physical volume evaluated in the scaling region [DiCa, ] turned out to be larger than in Wilson’s regularization.
@@
In conclusion, in four dimensions there is agreement between half-compact and compact regularizations for Abelian theories coupled to matter fields which however turn out to be both trivial, and between compact and non-compact regularizations of non-Abelian theories restricted to SU(2) and in the absence of matter fields.
@@
The paper is organized in the following way. In Section II we derive the definition of gauge transformations for Abelian fields and the relative Lagrangian including interactions with matter fields both in Cartesian and in polar representation, the latter being more convenient for a comparison with Wilson’s theory. Such results can be straightforwardly obtained from the general definition for non-Abelian theories [Palu1, ] but they have not been spelled out before. In Section III we derive the U(1) modified Lagrangian. In Section IV we report for the convenience of the reader the known results about models with gauge fields in the irreducible representation of SU(2) and construct the corresponding modified Lagrangians. All the above results are valid in four-dimensional Euclidean space. In Section V we briefly indicate how they should be modified to get the three-dimensional case. In Section VI we give our conclusions. We set everywhere $\hbar=c=1$.
II Non-compact U(1) gauge theories on a lattice
@@
We will derive the formulation of non-compact U(1) gauge theories on a lattice in four dimensions from the GL$(N_{c})$ ones [Palu1, ], which is straightforward but it has not been spelled out so far. We do it for a twofold motivation that it will be per sé interesting if, unlike the compact and half-compact regularization, it will turn out to be not trivial, and because it will serve to construct the three dimensional case in Section V.
@@ II.1 Gauge fields transformations
where $\mu$ in the argument of $g^{*}(x+\mu)$ represents the unit vector in the $\mu$-direction and $\theta(x)$ is the parameter of the transformation. For small $\theta(x)$ the gauge fields transform can be obtained from those of SU(2) [Becc1, ]
@@
II.2 The non-compact gauge fields Lagrangian
@@
It gives to the auxiliary field $W_{\mu}$ the mass $4\gamma^{2}/a^{2}$ which will play the role of an inverse expansion parameter. Moreover it damps all the gauge fields, including their zero-modes, so making unnecessary a gauge fixingaaaIt must be noticed that the potential $\mathcal{L}_{\rm m}$ has two minima: for $A_{\mu}=0$ $W_{\mu}=\frac{1\mp 1}{a}\,$ which should both be accounted for. The problem disappears using polar fields, as noted for SU(2) in Ref. [Becc2, ].
@@ II.3 Matter fields Lagrangians
This Lagrangian is usually called naive because it is plagued by the so called doubling problem, which has been tackled in various ways [Mont, ; Kaplan, ]. We will show in Sec. III how it can be handled in the non-compact formulation.
II.4 Polar fields
We derive the explicit expression of the non-compact Abelian Lagrangian with polar fields, which makes the comparison with Wilson’s theory easier.
@@ II.4 Polar fields
The potential $\mathcal{L}_{\rm m}$ is already written in Eq.(14) while the non-compact Yang-Mills Lagrangian density (12) in polar form is
@@
and, taking into account the symmetry of $\rho_{P\mu\nu}$ under $\mu\leftrightarrow\nu$
@@
is the $\rho_{\mu}$ Lagrangian. Notice that while $\mathcal{L}_{\rm YM}$ is local in Cartesian coordinates, it has been separated in non-local polar terms (non-polynomial when expressed in terms of Cartesian fields).
@@
If we want to stay strictly local this is not possible because $\mathcal{P}$ must be a polynomial in $\rho_{\mu}^{2}$. However, if we set
@@
in the approximation we will use $\mathcal{P}+\mathcal{L}_{\rho}+\mathcal{L}_{l}=0$, and thus:
@@
III Modified U(1) Lagrangians from the non-compact one
@@
where $\lambda$ is an arbitrary parameter with the dimension of a length. The potential term $\mathcal{L}_{\rm m}$ generates in such a case the $\delta$-function $\delta(\rho_{\mu}(x)-1/a)$ which eliminates the auxiliary field $\rho_{\mu}$ from the partition function.
But, based on the results [DiCa, ], which showed that the non-compact regularization generates for SU(2) a physical space larger than the compact one, we evaluate also for U(1) the contribution arising from the auxiliary field for small lattice spacing, and therefore we assume $\eta=0$.
@@ III.1 Modified compact Lagrangian
which shows that $\gamma$ (the mass of $s_{\mu}$-field, as it can be seen by the rescaling $s_{\mu}\rightarrow\gamma s_{\mu}$) plays the role of an inverse coupling constant. Therefore
@@
$\mathcal{L}^{\rm mod}_{\rm G}$ is the modified gauge fields Lagrangian, which for $\gamma\to\infty$ coincides with that of Wilson. For a first investigation one might just evaluate this contribution to the Wilson action at the phase transition.
Let us go to scalar fields, whose covariant derivative (omitting the flavour index) is
@@
$\mathcal{L}^{\rm mod}=\mathcal{L}_{\rm G}^{\rm W}+\mathcal{L}_{\phi}^{\rm W}-\frac{a^{4}}{2\gamma^{2}}\,\sum_{\mu}\left(\delta\mathcal{L}^{\phi}_{\mu}+\beta\,\sum_{\nu}U_{P\mu\nu}\right)^{2}\,.$ (50)
@@
Performing a by part integration in the action and expanding the $\rho$-field, we get to first order in the lattice spacing
@@
In order to see how can we account for the doubling problem in the present formalism we observe that we can perform the reverse transformation from compact to non-compact form by multiplying the Wilson variables $U_{\mu}$ by $a\rho_{\mu}$. Therefore, in all the cases in which the doubling problem is solved by addition of an appropriate gauge invariant term to the Lagrangian, one must make in such a term the replacement $U_{\mu}\,\to\,a\,\rho_{\mu}\,U_{\mu}$ and expand $\rho_{\mu}$ according to Eq. (39).
The final correction to $\mathcal{L}^{\rm W}_{\psi}$ has the same form as that of the scalar fields, and it can be handled the same way which yields, however, a quartic term in the fermion fields. Usually it is preferable instead to first integrate over the fermion Grassmann variables. We notice in this connection that the correction term might make easier the inversion of the fermion matrix in the chiral limit even though, at variance with the linearized gauge Nambu–Jona-Lasinio model [Kogu2, ], the help would not come from a mass-like term but from the fermion current.

### v2

# Modified Abelian and SU(2) Wilson theories on a lattice from a noncompact regularization

@@ Abstract
Multiflavor gauge theories of matter systems on a three-dimensional lattice have recently been widely investigated especially in connection with a possible symmetry enlargement at a continuous phase transition. Abelian models were studied both with compact gauge fields and in a mixed formulation in which the coupling with matter fields is in compact form while the gauge fields Lagrangian is written in terms of noncompact gauge fields, getting quite different results. Such a mixed formulation is not permissible for non-Abelian theories, for which however there exists an entirely noncompact formulation (in which exact gauge invariance is enforced by help of auxiliary fields), which for SU(2) was shown to yield in the scaling window a larger physical volume than Wilson’s one. The corresponding U(1) noncompact regularization is derived in the present work. In both Abelian and SU(2) cases there is only one auxiliary field that for a large mass has a linear coupling with the other fields and it can be integrated out yielding a negative definite correction. The coupling with the auxiliary field might make the inversion of the fermion matrix easier.
@@ I Introduction
The initial motivation of the present study was the recent interest in multiflavor gauge theories in three space-time dimensions [Peli1, ; Bona1, ; higher-charge, ; Bona2, ], but we think that the results have a wider perspective. Multiflavor gauge theories appear in the investigation of condensed matter [Klei, ] and critical phenomena, especially in connection with a possible symmetry enlargement at a second order phase transition [Sach, ]. The relative nonperturbative investigations [Peli1, ; Bona1, ; higher-charge, ; Bona2, ] have been mostly performed by lattice Monte Carlo simulations and in several Abelian models the phase diagrams turned out to depend on the way, compact or noncompact, the gauge fields are defined on a lattice.
@@
Later on a new regularization of GL($N_{c}$) theories on a lattice was defined [Palu1, ], in which the gauge fields are maintained in the algebra of the group, namely they are noncompact, and in order to enforce gauge invariance auxiliary fields are introduced which decouple in the continuum limit. Since also the auxiliary fields are noncompact, such a regularization can be called noncompact. For $N_{c}=2$ the irreducible representation has an SU(2) invariance [Becc1, ], while the general definition for SU($N_{c}>2$) symmetry has been given in a subsequent work [Palu2, ]. The construction of such noncompact regularization was done with the idea of being closer to the continuum at finite lattice spacing, an expectation supported by a Monte Carlo simulation in four dimensions which showed that in the scaling window it generates for SU(2) a physical space larger than the Wilson’s one [DiCa, ].
@@
In the study of Abelian theories another regularization [Bona1, ] is used, in which the gauge fields Lagrangian is written as a discretization of the continuous one, namely it is noncompact, while the coupling to matter fields is written in terms of Wilson variables. For this reason, and to distinguish it from the entirely noncompact regularization [Palu1, ] that we will use, we will call it half-compact, even though in the literature it is generally referred to as noncompact, hoping that this will not generate any confusion.
@@
Compact and half-compact regularizations in three dimensions give quite different phase diagrams. In the first case the flavor symmetry is broken while the gauge fields are in the confined phase; in the second case there are three phases: the Higgs and molecular phases, in which the flavor symmetry is broken and the electromagnetic correlations are gapped and ungapped, respectively, and the Coulomb phase in which the flavor symmetry is unbroken and the electromagnetic correlations are long range. In view of such striking difference it seems worthwhile to compare them with the entirely noncompact one [Palu1, ] that might also introduce further physical behaviors.
@@
The half-compact regularization cannot be used for non-Abelian theories, because it breaks the gauge invariance of gauge fields self-interactions and therefore a comparison between compact and half-compact was restricted to the Abelian case. Also for non-Abelian theories, however, a comparison can be made with the noncompact regularization, for which there is the additional expectation of a faster convergence to the continuum.
@@
As a first approach instead of studying entirely noncompact models, one might investigate the corrections they give to the Wilson Lagrangian at small lattice spacing, in analogy to the modifications obtained by using the renormalization group transformations [Wils2, ] or adding some extra terms chosen to cancel the leading lattice artifacts [Syma, ]. These methods will be discussed in Sec. V. It is the purpose of the present work to derive such corrections.
@@
SU(2) and U(1) theories, which are the ones investigated in the quoted multiflavor models [Peli1, ; Bona1, ; higher-charge, ; Bona2, ] have in their noncompact form one feature in common: they have only one auxiliary field that in polar variables is a gauge invariant function of the sites. For large mass this field has linear couplings with all the other (gauge and matter) fields. The resulting Lagrangian is analogous to the gauged Nambu–Jona-Lasinio model [Kogu2, ] after linearization: the auxiliary field can then be integrated out leaving a Lagrangian expressed in terms of Wilson variables and matter fields only. Of course, in the presence of fermion fields the integration over the auxiliary field has to be performed after that on the Grassmann variables, but in analogy with the gauged Nambu–Jona-Lasinio model, the presence of the auxiliary field might make easier the inversion of the fermion matrix (but see later on this point).
@@
The half-compact and noncompact regularizations differ from each other also in another point: only the first one needs a gauge fixing to get rid of the pure gauge degrees of freedom that have an infinite volume in the partition function. Some gauge fixings, however, are not compatible with some boundary conditions (see for instance Ref. [Bona3, ]),while the choice of boundary conditions may have physical consequences, a point illustrated by a few examples in Appendix IX. In the noncompact regularization [Palu1, ] instead no gauge fixing is necessary because a gauge invariant potential damps all the gauge fields, physical and auxiliary, so that all boundary conditions can be handled on the same footing.
@@
Abelian theories have been compared in the compact and half-compact forms. It was soon established that the first one does not yield a second order phase transition in the bare coupling constant and therefore it cannot define a continuum theory [Kogu1, ]. Several investigations have been performed afterwards using the half-compact formulation. After the result that at small coupling the theory is trivial also in this case, it was searched whether a continuum theory could be defined at finite coupling [Azco1, ], until it was established that models of Abelian gauge fields coupled to scalars [Baig, ] or fermions with four flavors in the chiral limit [Kogu2, ] are indeed trivial. Later on, however, evidence was reported that in a large-$N$ expansion an ultraviolet fixed point exists in a theory of gauge fields regularized in the continuum and coupled to both fermions and scalars [Liti, ].
@@
Non-Abelian theories instead have been studied in compact and noncompact form, the latter one however restricted to SU(2) and only in the absence of matter fields. First, it was established that the noncompact formulation gives an evidence for confinement [Poli1, ] comparable to that of Wilson. This was a prerequisite, because noncompact gauge fields with a simple discretization of the continuum theory [Patr, ] or a nonrenormalizable gauge fixing [Poli2, ] gave a vanishing string tension. Next, a perturbative study of the scaling properties determined the renormalization scale parameter and showed that the theory has the same fixed point as Wilson’s one [Becc2, ]. This result was confirmed by a calculation [Bora, ] in the Hamiltonian formalism [Dick, ] using Lüscher’s effective Hamiltonian [Lusc, ]. Lastly, as already said, the physical volume evaluated in the scaling region [DiCa, ] turned out to be larger than in Wilson’s regularization.
@@
In conclusion, in four dimensions there is agreement between half-compact and compact regularizations for Abelian theories coupled to matter fields that however turn out to be both trivial (but it should be reexamined in the presence of coupling to both fermions and scalars) and between compact and noncompact regularizations of non-Abelian theories restricted to SU(2) and in the absence of matter fields.
@@
The paper is organized in the following way. In Sec. II we derive the definition of gauge transformations for Abelian fields and the relative Lagrangian including interactions with matter fields both in Cartesian and in polar representation, the latter being more convenient for a comparison with Wilson’s theory. Such results can be straightforwardly obtained from the general definition for non-Abelian theories [Palu1, ] but they have not been spelled out before. In Sec. III we derive the U(1) modified Lagrangian. In Sec. IV we report for the convenience of the reader the known results about models with gauge fields in the irreducible representation of SU(2) and construct the corresponding modified Lagrangians. In Sec. V we discuss some previous improvements of the Wilson Lagrangian.
All the above results are valid in four-dimensional Euclidean space. In Sec. VI we briefly indicate how they should be modified to get the three-dimensional case. In Sec. VII we give our conclusions. We set everywhere $\hbar=c=1$.
II Noncompact U(1) gauge theories on a lattice
@@
We will derive the formulation of noncompact U(1) gauge theories on a lattice in four dimensions from the GL$(N_{c})$ ones [Palu1, ], which is straightforward but it has not been spelled out so far. We do it for a twofold motivation: the possibility that unlike the compact and half-compact regularization it will turn out not trivial (in the absence of coupling with both fermions and scalars), and because it will serve to construct the three-dimensional case in Sec. VI.
@@ II.1 Gauge fields transformations
where $\mu$ in the argument of $g^{*}(x+\mu)$ represents the unit vector in the $\mu$ direction and $\theta(x)$ is the parameter of the transformation. For small $\theta(x)$ the gauge fields transformation can be obtained from those of SU(2) [Becc1, ]
@@
II.2 The noncompact gauge fields Lagrangian
@@
It gives to the auxiliary field $W_{\mu}$ the mass $4\gamma^{2}/a^{2}$ which will play the role of an inverse expansion parameter. Moreover it damps all the gauge fields, including their zero modes, so making unnecessary a gauge fixingaaaIt must be noticed that the potential $\mathcal{L}_{\rm m}$ has two minima: for $A_{\mu}=0$ $W_{\mu}=\frac{1\mp 1}{a}\,$ which should both be accounted for. The problem disappears using polar fields, as noted for SU(2) in Ref. [Becc2, ]..
@@
This Lagrangian is usually called naive because it is plagued by the so-called doubling problem [Mont, ; Kaplan, ; PaluNC, ] that has been tackled in various ways and it can be handled in the noncompact formulation as indicated in Sec. III. Reference [PaluNC, ] contains a second order formulation of free chiral fermions on a lattice. Addition of gauge interactions should not give any problem, but a difficulty can be encountered to project out in a gauge-invariant way zero modes of auxiliary scalars in the spinor representation.
II.3 Polar fields
We derive the explicit expression of the noncompact Abelian Lagrangian with polar fields, which makes the comparison with Wilson’s theory easier.
@@ II.3 Polar fields
The potential $\mathcal{L}_{\rm m}$ is already written in Eq.(14) while the noncompact Yang-Mills Lagrangian density (12) in polar form is
@@
and, having taken into account the symmetry of $\rho_{P\mu\nu}$ under $\mu\leftrightarrow\nu$,
@@
is the $\rho_{\mu}$ Lagrangian. Notice that while $\mathcal{L}_{\rm YM}$ is local in Cartesian coordinates, it has been separated in nonlocal polar terms (nonpolynomial when expressed in terms of Cartesian fields).
@@
If we want to stay strictly local, then this is not possible because $\mathcal{P}$ must be a polynomial in $\rho_{\mu}^{2}$. However, if we set
@@
in the approximation we will use $\mathcal{P}+\mathcal{L}_{\rho}+\mathcal{L}_{l}=0$, and thus
@@
III Modified U(1) Lagrangians from the noncompact one
@@
where $\lambda$ is an arbitrary parameter with the dimension of a length. The potential term $\mathcal{L}_{\rm m}$ generates in such a case the $\delta$ function $\delta(\rho_{\mu}(x)-1/a)$ that eliminates the auxiliary field $\rho_{\mu}$ from the partition function.
But, based on the results [DiCa, ], which showed that the noncompact regularization generates for SU(2) a physical space larger than the compact one, we evaluate also for U(1) the contribution arising from the auxiliary field for small lattice spacing, and therefore we assume $\eta=0$.
@@ III.1 Modified compact Lagrangian
which shows that $\gamma$ (the mass of the $s_{\mu}$ field, as it can be seen by the rescaling $s_{\mu}\rightarrow\gamma s_{\mu}$) plays the role of an inverse coupling constant. Therefore
@@
$\mathcal{L}^{\rm mod}_{\rm G}$ is the modified gauge fields Lagrangian, which for $\gamma\to\infty$ coincides with that of Wilson while at finite $\gamma$ contains a negative definite contribution. For a first investigation one might just evaluate this contribution to the Wilson action at the phase transition.
Let us go to scalar fields, whose covariant derivative (omitting the flavor index) is
@@
$\mathcal{L}^{\rm mod}=\mathcal{L}_{\rm G}^{\rm W}+\mathcal{L}_{\phi}^{\rm W}-\frac{a^{4}}{2\gamma^{2}}\,\sum_{\mu}\left(\delta\mathcal{L}^{\phi}_{\mu}+\beta\,\sum_{\nu}U_{P\mu\nu}\right)^{2}$ (50)
that also contains a negative definite contribution.
@@
Performing a by part integration in the action and expanding the $\rho$ field, we get to first order in the lattice spacing
@@
In order to see how can we account for the doubling problem in the present formalism we observe that we can perform the reverse transformation from compact to noncompact form by multiplying the Wilson variables $U_{\mu}$ by $a\rho_{\mu}$. Therefore, in all the cases in which the doubling problem is solved by addition of an appropriate gauge invariant term to the Lagrangian, one must make in such a term the replacement $U_{\mu}\,\to\,a\,\rho_{\mu}\,U_{\mu}$ and expand $\rho_{\mu}$ according to Eq. (39).
The final correction to $\mathcal{L}^{\rm W}_{\psi}$ has the same form as that of the scalar fields, and it can be handled the same way that yields, however, a quartic term in the fermion fields. Usually it is preferable instead to first integrate over the fermion Grassmann variables. We notice in this connection that the correction term might make easier the inversion of the fermion matrix in the chiral limit even though, at variance with the linearized gauge Nambu–Jona-Lasinio model [Kogu2, ], the help would not come from a masslike term but from the fermion current.